# The Cone Restriction: An old Approach Revisited

### v5

@@ 2.2 Wave Packet Decomposition
[…]
@@
here ${\mathrm{D}\mathrm{ist}}_{H}$ denotes the Hausdorff distance. For a $r$-tube $T^{l}_{\theta,v}$, if there are multiple $\left({\zeta}_{0},v_{0}\right)$ satisfying the above conditions, then we assign $T^{l}_{\theta,v}$ to only one ${\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)$. Also, we also define
@@
One immediately concludes that
${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B\left(v_{0},r^{\frac{1}{2}+2\delta}\right),\ \ \forall\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}},\quad g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.2)
Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
@@
From Lemma 2.5 and 2.6, one has
@@
[…]
Tangential and transverse wave packets Here we recall some definitions and properties about the tangential and transverse wave packets from [6, 15].
@@
From the above definition, if $\mathrm{Z}:=Z\left(P_{1},\dots,P_{n-m}\right)$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$, in other words, $\mathrm{Z}$ has a tangent plane $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
@@
Let $\boldsymbol{\mathrm{Z}}$ be a transverse complete intersection of dimension $m$. For a fixed ball $B\left(y,r\right)$, a $r$-tube $T^{l}_{\theta,v}$ with $T^{l}_{\theta,v}\cap B\left(y,r\right)\cap\boldsymbol{\mathrm{Z}}\neq\emptyset$ is said to be tangent to $\mathrm{Z}$ in $B\left(y,r\right)$ if $T^{l}_{\theta,v}\cap B\left(y,2r\right)\subseteq N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)$ and
$\angle\left(L\left(\theta\right),T_{x}\boldsymbol{\mathrm{Z}}\right)\lesssim r^{-1/2+{\delta}_{m}},\ \ \forall x\in\mathrm{Z}\cap B\left(y,r\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(T^{l}_{\theta,v})\right)$
Otherwise, we say that $T^{l}_{\theta,v}$ is transverse to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$. We use $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the $r$-tube tangent to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$.
In Section 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
@@
$Z_{>\alpha}:=\left\{x\in Z:\angle\left(l,T_{x}\boldsymbol{\mathrm{Z}}\right)>\alpha\right\}$
@@
2.3 Some Kakeya-type geometric results
First, we introduce some definitions about grains and multigrain from Section 3 of [12].
Definition 2.11 ([12], Definition 3.3).
A grain is defined to be a pair $\left(Z,B\left(y,r\right)\right)$ where $Z$ is a transverse complete intersection and $B_{r}$ is a ball of some radius $r>0$. The dimension of a grain $\left(Z,B\left(y,r\right)\right)$ is that of the variety $Z$, while its degree is that of $Z$ and its scale is $r$.
Definition 2.12 ([12], Definition 3.4).
Let $S=\left(Z,B\left(y,r\right)\right)$ be a grain of dimension $m$. A function $f$ with $\mathrm{supp}\ f\subseteq 2{\overline{B}}^{n-1}\backslash B^{n-1}$ is said to be tangent to $S$ if it is concentrated on scale $r$ wave packets belong to the collection:
Definition 2.13 ([12], Definition 3.5).
A multigrain is a tuple of grains
satisfying
${\mathrm{dim}Z_{i}\ }=i,\quad\forall i$
$B(y_{i},r_{i})\cap B(y_{i+1},r_{i+1})\neq\emptyset,\quad\forall i$
$0<r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
The parameter $\left(t-s+1\right)$ is called the multigrain level. The degree of $\overrightarrow{S}$ is defined as $\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\$}. The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=\left(r_{s},r_{s+1},\dots,r_{t}\right)$. Finally, we write
The restriction conjecture is associated with a Kakeya-type problem concerning the continuum incidence theory of tubes. Specifically, given a multigrain ${\overrightarrow{S}}_{s,t}=\left(S_{s},\dots,S_{t}\right)$ with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$, if $\mathcal{L}$ is a set of $r^{-1/2}_{t}$-direction separated rays that satisfies
$L\cap B\left(y_{i},r_{i}\right)\cap N_{{\rho}_{j}}Z_{i}\neq\emptyset$
$L\cap B\left(y_{i},2r_{i}\right)\subseteq N_{{\rho}_{j}}Z_{i}\cap B\left(y_{i},2r_{i}\right)$
for any $L\in\mathcal{L}$, then what is the maximum possible value of $\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
$\angle\left(L_{1},L_{2}\right)\gtrsim r^{-\frac{1}{2}},\ \ \mathrm{if}\ L_{1},L_{2}\in\mathcal{L}\boldsymbol{\ }\mathrm{and}\boldsymbol{\ }L_{1}\neq L_{2}$
[…]
Theorem 2.14 (Polynomial Wolff Axiom).
Let $r\geq 1$. Let $Z\subset{\mathbb{R}}^{n}$ be an algebraic variety of dimension $m$ and degree at most $d$. If $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$satisfies: $\forall T\in\mathcal{T}$,
$T\cap B_{r}\cap N_{r^{1/2+{\delta}_{m}}}Z\neq\emptyset\$
$T\cap 2B_{r}\subseteq N_{r^{1/2+{\delta}_{m}}}Z$
then
$\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
[…]
Definition 2.15 (Nested Tubes).
and
[…]
Finally, we set
One can observe that the second condition in Definition 2.15 is weaker than Lemma 2.5. However, this is good enough for our arguments. In [12, 8, 10] a nontrivial extension of Theorem 2.14, which is a corollary to Lemma 2.11 in [24], was used respectively to establish a new result on the Bochner-Riesz problem, Stein’s square function, and the Fourier restriction for paraboloid. The result is shown below:
Lemma 2.16 ([12], Lemma 3.7, Nested polynomial Wolff axioms).
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
A direct corollary of this lemma is
Corollary 2.17.
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
Proof.
In [12, 8, 10] $S_{m}$ is traced back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus Lemma 2.16 gives
$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
Since the tubes have $\sim R^{\left(n-1\right)/2}$ directions in the case of paraboloids, this bound yields a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$. However, in the case of cone, this does not guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$. This is the why we need Corllary 2.17
2.4 $k$-broad Norms
[…]
$L\left(\tau\right):=\left\{L\left(\theta\right):\theta\subset\tau\right\}$
@@ 2.4 $k$-broad Norms
Cover $B_{R}$ by finitely overlapping $B_{K^{2}}$. Fix a parameter $1<A\lesssim K^{\epsilon}$. For each $B_{K^{2}}$, define
@@
${\mu}_{Ef}\left(B_{K^{2}}\right):={\mathop{\mathrm{min}}_{V_{1},\dots,\ V_{A}\ \left(k-1\right)\mathrm{-}\mathrm{subspace\ of}\ {\mathbb{R}}^{n}}\left(\mathop{\mathrm{max}}_{\tau:\angle\left(L\left(\tau\right),V_{a}\right)\gtrsim K^{-2},\forall 1\leq a\leq A}\int_{B_{K^{2}}}{{\left|Ef_{\tau}\right|}^{p}}\right)\ }$
here
$\angle\left(L\left(\tau\right),V_{a}\right):={\mathop{\mathrm{min}}_{v\in G\left(\tau\right)}\angle\left(v,V_{a}\right)\ }$
Compared with the paraboloid case [6], the angel condition is modified to $\angle\left(G\left(\tau\right),V_{a}\right)\gtrsim K^{-2}$, due to some geometric feature of the cone. See [15] for details. For $U\subseteq{\mathbb{R}}^{n}$ the $k$-board norm over $U$ is defined as follows
@@
[…]
Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about ${\left\|.\right\|}_{BL^{p}_{k,A}}$.
@@
Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If $\mathrm{deg}\boldsymbol{Z}\lesssim 1\$} and $dim\boldsymbol{Z}\$} $\leq k-1$, then
@@
These inequalities will be used frequently in Section 3.
Now suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(0,r\right)\right]$, where $\boldsymbol{\mathrm{Z}}$ is a $m$-dimensional transverse complete intersection. Fix a ball $B\left(y,\rho\right)\subseteq B\left(0,r\right)$. To apply the induction on scale, we expect to find a decomposition
$\tilde{g}=\sum_{b\in\mathcal{B}}{{\tilde{g}}_{b}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.4)
where
${\tilde{g}}_{b}:=\sum_{\left(\zeta,w,L\right)\in\mathbb{T}\left[{\boldsymbol{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]}{{\tilde{g}}^{L}_{\zeta,w}},\ \ {\boldsymbol{\mathrm{Z}}}_{b}:=\left\{z+b:z\in\boldsymbol{\mathrm{Z}}\right\}$
Roughly speaking, for each $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$, $N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)\cap B\left(y,\rho\right)$ is a ‘slice’ of $N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)\cap B\left(y,\rho\right)$. Let $g_{b}\left(\xi\right):=e^{-i{\psi}_{y}\left(\xi\right)}{\tilde{g}}_{b}\left(\xi\right)$, then
$Eg_{b}\left(x\right)\approx{\chi}_{N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)}\left(x\right)Eg\left(x\right)$ (2.5)
If $\boldsymbol{\mathrm{Z}}$ is a plane, then such a decomposition exists due to Lemma 2.6. However, this question will be tricky if ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular. A rigorous version of (2.4) and (2.5) is given by the following statement. See [6, 15, 4] for details.
Lemma 2.22.
[…]
${\left\|Eg\right\|}^{p}_{BL^{p}_{k,A}\left(B\left(y,\rho\right)\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in\mathcal{B}}{{\left\|Eg^{\prime}_{b}\right\|}^{p}_{BL^{p}_{k,A}\left(B_{\rho}\cap N_{{\rho}^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}+b\right)\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
and
[…]
where
$g^{\prime}:=\sum_{\left(\theta,v,l\right)\in{\mathcal{T}}^{\prime}}{g^{l}_{\theta,v}}$
[…]
Reduction to Broad Estimates The main result of this paper is
Theorem 2.23.
Let $2\leq k\leq n-2$ and
$p>p_{n}\left(k\right):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$ (2.6)
Then there exists an integer $A$ such that
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}{\lesssim}_{p,\varepsilon}R^{\varepsilon}{\left\|f\right\|}^{2/p}_{2}{\left\|f\right\|}^{1-2/p}_{\infty}$ (2.7)
Theorem 2.23, the $k$-broad estimate, implies Theorem 1.5, the linear cone restriction estimate. In particular, it follows from
@@
For any $R>1$, the estimate (1.3) holds whenever the estimate (2.7) holds, for $p$ in the range
$p\left(k,n\right)<p\leq\frac{2n}{n-2},p\left(k,n\right):=\left\{\begin{array}[]{c}2+\frac{2}{n-2},\ if\ 2\leq k\leq 3\\ 2+\frac{4}{2n-k-1},\ if\ k>3\end{array}\right$. (2.8)
We choose a value of $k$ to optimizes the range of $p$ in (1.3). By applying the arguments in [12] to (2.6), one has
$2+\frac{12}{4n-5+2\left(k-1\right){\left(\frac{k}{n-1}\right)}^{1/2}}\leq p_{n}\left(k\right)\leq 2+\frac{12}{4n-5+2\left(k-1\right){\left(\frac{2n-1}{2k+1}\right)}^{1/2}\frac{k}{n-1}}$
By writing $k=vn+O\left(1\right)$, one has
$p_{n}\left(k\right)=2+\frac{6}{2+v^{3/2}}n^{-1}+O\left(n^{-2}\right)$
From (2.8),
$p\left(k,n\right)=2+\frac{4}{2-v}n^{-1}+O\left(n^{-2}\right)$ (2.9)
To obtain the best possible estimate in Theorem 1.5, $v$ should satisfy
$\frac{6}{2+v^{3/2}}=\frac{4}{2-v}$
Hence, we choose
$v^{\frac{1}{2}}={\left(\frac{3}{8}+\frac{1}{2\sqrt{2}}\right)}^{1/3}+{\left(\frac{3}{8}-\frac{1}{2\sqrt{2}}\right)}^{1/3}-\frac{1}{2}=0.67765\dots$
Plugging this back into (2.9) yields the linear restriction estimate (1.3) holds for
$p>2+\lambda n^{-1}+O\left(n^{-2}\right)$
where
$\lambda=\frac{4}{2-v}=2.59607\dots$
This gives Theorem 1.5.
2.5 Transverse Equidistribution Estimates
By using a version of Heisenberg uncertainty principle, Guth [6] established the transverse equidistribution estimates for paraboloids. However, due to the thin thickness of tubes in the case of cone, $Eg$ may be concentrated on an extremely thin layer of $N_{R^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)$, and thus this property may fail to hold.
Figure 3: A scenario that suggests the transverse equidistribution fails.
For example, let $\boldsymbol{\mathrm{Z}}$ be a $\left(n-1\right)$-plane, and $B\left(y,r\right)\subseteq B\left(0,R\right)$ for some $y\in\boldsymbol{\mathrm{Z}}$ and $r\sim R^{2/3}$. Let $T^{l}_{\theta,v}$ be a $R$-tube with $T^{l}_{\theta,v}\subseteq N_{R^{1/2+{\delta}_{n-1}}}\left(\boldsymbol{\mathrm{Z}}\right)\cap B\left(0,R\right)$, $L\left(\theta\right)\parallel\boldsymbol{\mathrm{Z}}$, and $M\left(\theta\right)\bot\boldsymbol{\mathrm{Z}}$. Suppose $g$ is concentrated on $T^{l}_{\theta,v}$. Since $r^{1/2+{\delta}_{n-1}}\gg R^{\delta}$, there exist $b\in B\left(0,R^{1/2+{\delta}_{m}}\right)$ such that $T^{l}_{\theta,v}\subseteq N_{r^{1/2+{\delta}_{n-1}}}\left(\boldsymbol{\mathrm{Z}}\mathrm{+b}\right)\cap B\left(0,R\right)$. Therefore,
${\left\|Eg_{b}\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}={\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}^{2}_{2}$
Since $r/R\sim R^{-1/3}$, one has
${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$
If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then Lemma 2.4 gives
${\left\|g_{b}\right\|}^{2}_{2}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|g\right\|}^{2}_{2}$
which suggests the transverse equidistribution estimates fail.
[…]

### v7

@@ 2.1. Wave Packet Decomposition
Roughly speaking, Lemma 2.6 asserts that $(g_{\theta,v}^{l})^{\widetilde{}}$ is concentrated on scale $\rho$ wave packets that are parallel to $T_{\theta,v}^{l}$ and contained in $c\,T_{\theta,v}^{l}$ for some constant $c$.
2.4. Medium Tubes.
Next, we group $\mathbb{T}[r]$ and $\mathbb{T}[\rho]$ into subcollections. For a $\rho^{-1/2}$-sector $\zeta_{0}$ and $v_{0}\in r^{1/2+2\delta}\mathbb{Z}^{n-1}\cap B(0,\rho)$, let $\mathbb{T}_{\zeta_{0},v_{0}}(y)$ denote the collection of $r$-tubes $T_{\theta,v}^{l}$ such that
@@
here $\mathrm{Dist}_{H}$ denotes the Hausdorff distance. If an $r$-tube $T_{\theta,v}^{l}$ satisfies the above conditions for multiple pairs $(\zeta_{0},v_{0})$, we assign $T_{\theta,v}^{l}$ to only one $\mathbb{T}_{\zeta_{0},v_{0}}(y)$. We also define
@@
We immediately conclude that for all $(\zeta,w,L)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$
${\mathrm{Dist}}_{H}({\zeta}_{0},\zeta)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B(v_{0},r^{\frac{1}{2}+2\delta})$
and
$g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$
Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, Lemma 2.1 implies
@@
It follows from Lemmas 2.5 and 2.6 that
@@
2.5. Tangential and Transverse Wave Packets.
We recall several definitions and properties of tangential and transverse wave packets from [6, 15].
@@
From the above definition, if $\mathrm{Z}:=Z(P_{1},\dots,P_{n-m})$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$. In particular, $\mathrm{Z}$ admits a tangent space $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
@@
Let $\bm{\mathrm{Z}}$ be a transverse complete intersection of dimension $m$. For a fixed ball $B\left(y,r\right)$, a $r$-tube $T^{l}_{\theta,v}$ with $T^{l}_{\theta,v}\cap B\left(y,r\right)\cap\bm{\mathrm{Z}}\neq\emptyset$ is said to be tangent to $\mathrm{Z}$ in $B\left(y,r\right)$ if $T^{l}_{\theta,v}\cap B\left(y,2r\right)\subseteq N_{r^{1/2+{\delta}_{m}}}\left(\bm{\mathrm{Z}}\right)$ and
$\angle\left(L\left(\theta\right),T_{x}\bm{\mathrm{Z}}\right)\lesssim r^{-1/2+{\delta}_{m}},\ \ \forall x\in\mathrm{Z}\cap B\left(y,r\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(T^{l}_{\theta,v})\right)$
Otherwise, we say that $T^{l}_{\theta,v}$ is transverse to $\bm{\mathrm{Z}}$ in $B\left(y,r\right)$. We use $\mathbb{T}\left[\bm{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the set of $r$-tubes tangent to $\bm{\mathrm{Z}}$ in $B\left(y,r\right)$.
In Section 3, we require the following result, which can be viewed as a continuum analog of Bézout’s theorem, to control the transverse intersections between a tube and an algebraic variety.
@@
$Z_{>\alpha}:=\left\{x\in Z:\angle\left(l,T_{x}\bm{\mathrm{Z}}\right)>\alpha\right\}$
@@
2.6. Reduction to Broad Estimates
In [6], Guth decomposed $\|Ef\|_{L^{p}(B_{R})}$ into a broad part and a narrow part. The narrow part is essentially supported in a lower-dimensional subspace of $\mathbb{R}^{n}$. Using decoupling [2] together with induction on scales, he reduced the linear restriction estimates to estimating the broad part. Ou–Wang [15] followed the same strategy.
Fix a large constant $K$. Decompose $2\overline{B}^{\,n-1}\setminus B^{n-1}$ in frequency space into $K^{-1}$-sectors $\tau$. Given a Schwartz function $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$, write
$f=\sum_{\tau}f_{\tau},\qquad f_{\tau}:=f\,\psi_{\tau},$
@@ 2.6. Reduction to Broad Estimates
where $\{\psi_{\tau}\}_{\tau}$ is a smooth partition of unity subordinate to the cover $\{\tau\}$. Cover $B_{R}$ by finitely overlapping $B_{K^{2}}$. Fix a parameter $1<A\lesssim K^{\epsilon}$. For each $B_{K^{2}}$, define $\mu_{Ef}(B_{K^{2}})$ by
@@
$\mu_{Ef}(B_{K^{2}}):={\mathop{\mathrm{min}}_{V_{1},\dots,\ V_{A}\ \left(k-1\right)\mathrm{-}\mathrm{subspace\ of}\ {\mathbb{R}}^{n}}\left(\mathop{\mathrm{max}}_{\tau:\angle\left(L\left(\tau\right),V_{a}\right)\gtrsim K^{-2},\forall 1\leq a\leq A}\int_{B_{K^{2}}}{{\left|Ef_{\tau}\right|}^{p}}\right)\ }$
where
$\angle\left(L\left(\tau\right),V_{a}\right):={\mathop{\mathrm{min}}_{v\in L\left(\tau\right)}\angle\left(v,V_{a}\right)\ }$
For $U\subseteq{\mathbb{R}}^{n}$ the $k$-broad norm over $U$ is defined by
@@
Here we recall some basic properties about broad norm.
@@
Suppose $Ef$ is tangent to a transverse complete intersection $\bm{Z}$ in $B_{r}$. If $\mathrm{deg}\bm{Z}\lesssim 1\$} and $dim\bm{Z}\$} $\leq k-1$, then
@@
Broad estimates may imply linear estimates. More precisely,
@@
Let $n\geq 3$ and $2\leq q\leq p$. Suppose for some integer $A>0$ the estimate
$\|Ef\|_{BL_{k,A}^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$ (2.3)
holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$. If there exists an integer $k$ with $2\leq k\leq n$ and
$\left\{\begin{array}[]{rcl}q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
then the estimate (1.7) holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
Hence, together with Lemma 2.12, Proposition 1.7 reduces to
Proposition 2.16.
Suppose $n\geq 3$ and $2\leq k\leq n-2$. For $p=p(n,k)$, $q=q(n,k)$ and some integer $A\lesssim_{\varepsilon,n}1$, the estimate (2.3) holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
2.7. Transverse Equidistribution Estimates
Using a version of the Heisenberg uncertainty principle, Guth [6] established transverse equidistribution estimates for the paraboloid. However, in the cone setting the tubes are much thinner, and $Eg$ may be concentrated on an extremely thin layer of $N_{R^{1/2+\delta_{m}}}(\mathbf{Z})$, so this property may fail to hold.
To overcome this geometric obstruction, [15] decomposes $g$ into an essential part $g_{\mathrm{ess}}$ and a tail part $g_{\mathrm{tail}}$. Roughly speaking, $g_{\mathrm{ess}}$ is concentrated on wave packets $(\theta,v,l)$ for which $M(\theta)$ is not perpendicular to the tangent space of $\mathbf{Z}$ at some point $z\in\mathbf{Z}\cap N_{R^{1/2+\delta_{m}}}(T_{\theta,v}^{l})$; consequently, transverse equidistribution estimates hold for $g_{\mathrm{ess}}$. The tail term $g_{\mathrm{tail}}$ can be made negligible by choosing $A$ sufficiently large.
Let $\bm{\mathrm{Z}}$ be an $m$-dimensional transverse complete intersection with degree $\leq d$. Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{\mathrm{Z}}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{\mathrm{Z}},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$, where